Find the center and radius of the sphere given by the following equation.

x +y +z −4x+8y−12z+31=0
Answer #1

Given y + 3 - unt 89 - 122 + 3) = 0 6 - 4x) + ( + 8) + (2 - 123) +31 = 0 - 123+36-36) - (x - 4x + 4-4) + (y +49 +16-16) + (22